OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. ( 4 q ^-1 /z^ -3 )^ -2

OpenStudy (saifoo.khan):

\[\Large (\frac{4q^{-1}}{z^{-3}})^{-2}\]\[(\Large \frac{z^{-3}}{4q^{-1}})^2\]

OpenStudy (anonymous):

\[(\frac{4q^{-1}}{z^{-3}})^{-2}\]\[(\frac{1}{4q}\times z^3)^{-2}\]\[\frac{16q^2}{z^6}\]

OpenStudy (saifoo.khan):

kmousou's right.

OpenStudy (anonymous):

@bradley0516 Is it \[(4q)^-{1}\]or\[4q^-{1}\]?

OpenStudy (anonymous):

(4q ) -1 i think

OpenStudy (anonymous):

of courseit does say for positive exponent so it may be positive instead of negative

OpenStudy (anonymous):

If so, my first answer is correct, otherwise it is \[(\frac{4q^{-1}}{z^{-3}})^{-2}\]\[(\frac{4z^{3}}{q})^{-2}\]\[\frac{q^2}{16z^6}\]